OpenStudy (anonymous):

i will medal if its right @sleepyjess @horsegirl27 @MikeyMaximum please help

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

here

OpenStudy (alexandervonhumboldt2):

do yuo know distibutive property?

OpenStudy (alexandervonhumboldt2):

a*(b+c)=ab+ac

OpenStudy (sleepyjess):

Which do you think LuLu?

OpenStudy (alexandervonhumboldt2):

what you think?

OpenStudy (anonymous):

i think D maybe

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

B

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

C

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

7*3x=21x 7*11=77

OpenStudy (alexandervonhumboldt2):

21x-11x=10x

OpenStudy (anonymous):

ok i have more

OpenStudy (alexandervonhumboldt2):

10x+77 got!! yes tag us